# Functional differential equations driven by càdlàg rough paths

### v2

@@ 1. Introduction
where the driving signal $\mathbf{X}$ is a càdlàg $p$-rough path for $p\in(2,3)$, the initial condition $y$ is a given controlled path, and the coefficients $b,\sigma$ are non-anticipative functionals, mapping a controlled path to a controlled path. Assuming a quadratic growth and a Lipschitz type condition on the path-dependent coefficients $b,\sigma$, which both are formulated on the space of controlled paths, we establish the existence of a unique solution to the RFDE (2.2). To that end, we rely on the theory of càdlàg rough paths, as introduced by Friz, Shekhar and Zhang [FS17, FZ18], as well as Banach’s fixed point theorem. Moreover, we show that the solution map, also known as Itô–Lyons map, mapping the input (initial condition, coefficients, driving signal) of an RFDE to its solution, is locally Lipschitz continuous with respect to suitable distances on the associated spaces of coefficients, controlled paths and rough paths. Let us remark that the continuity of the Itô–Lyons map is one of the most fundamental insights of rough path theory, with many applications to stochastic differential equations, cf. e.g. [FH20].
@@
In the existing literature, there are several different approaches to deal with rough functional differential equations driven by continuous rough paths. Since the theory of (continuous) rough paths works nicely for infinite dimensional Banach spaces, RDEs with path-dependent coefficients can be treated as Banach space-valued RDEs, see e.g. [Bai14], which requires the coefficients to be Fréchet differentiable and, thus, excludes some interesting examples. Existence, uniqueness and stability results are established by Neuenkirch, Nourdin and Tindel [NNT08] for RDEs with constant delay. The existence of a solution is proven by Ananova [Ana23] for RDEs with path-dependent coefficients, which are assumed to be Dupire differentiable [Dup19], and by Aida [Aid24] for RDEs with coefficients containing path-dependent bounded variation. The latter two approaches rely on Schauder’s fixed point theorem. Another exemplary class of RFDEs are reflected rough differential equations, see e.g. [Aid15, DGHT19], which, in general, do not possess a unique solution, see [Gas21].
@@
Most applications of rough path theory to stochastic differential equations (SDEs) crucially rely on the construction of suitable (random) rough paths. To apply the developed theory on RFDEs to Itô SDEs with constant delays, we show that a càdlàg martingale together with its delayed version can be lifted to a random rough path in the spirit of stochastic Itô integration. The key challenge to obtain the ‘delayed’ rough path is that a martingale together with its delayed version is, in general, not a martingale itself, thus preventing the direct use of stochastic Itô integration. For related constructions of random rough paths above fractional Brownian motions we refer to [NNT08, TT12, BMCR14, CE20]. Consequently, one can apply the continuity of the Itô–Lyons map to derive pathwise stability results for stochastic differential equations with constant delay, which plays an important role in many applications, see e.g. [BDNHP18]. In particular, the map $y\mapsto Y$, mapping the initial condition $y$ to the associated solution $Y$ of an SDE with constant delay, is continuous on the space of controlled paths. This resolves an old observation, pointed out by Mohammed [Moh86], about the non-continuity of the flow of stochastic differential equations with delay, for which the initial condition is in fact an initial path.
@@
The organization of the paper is as follows. In Section 2 we provide existence, uniqueness and continuity results for rough functional differential equations. In Section 3 we prove that various classes of rough differential equations are covered by the presented results on rough functional differential equations. In Section 4 we establish the existence of the Itô rough path lift of delayed martingales and discuss applications to stochastic differential equations with delay. Appendix A contains some auxiliary estimates for rough integrals.
@@ 2.1. Essentials on rough path theory
The space of linear maps from $\mathbb{R}^{d}\to\mathbb{R}^{n}$ is denoted by $\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{n})$ and we write $C^{l}_{b}=C^{l}_{b}(\mathbb{R}^{n};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{n}))$ for the space of $l$-times differentiable (in the Fréchet sense) functions $f\colon\mathbb{R}^{n}\to\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{n})$ such that $f$ and all its derivatives up to order $l$ are continuous and bounded. We equip this space with the norm
@@
$X_{s,t}:=X_{t}-X_{s},\qquad\text{for}\quad(s,t)\in\Delta_{T}.$
@@
$\|\mathbf{X};\tilde{\mathbf{X}}\|_{p}:=\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[0,T]}\qquad\text{with}\qquad\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]}:=\|X-\tilde{X}\|_{p,[s,t]}+\|\mathbb{X}-\tilde{\mathbb{X}}\|_{\frac{p}{2},[s,t]},$
@@
$\|Y\|_{p}\leq C_{p}(\|Y^{\prime}\|_{\infty}\|X\|_{p}+\|R^{Y}\|_{\frac{p}{2}})\qquad\text{and}\qquad\|Y^{\prime}\|_{\infty}\leq|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p}.$
Given $X,\tilde{X}\in D^{p}$, we further introduce the standard ‘distance’
$\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p}:=\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[0,T]}$
@@
$\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[s,t]}:=|Y^{\prime}_{s}-\tilde{Y}^{\prime}_{s}|+\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[s,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[s,t]}$
for $(s,t)\in\Delta_{T}$, whenever $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X},(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}$. Note that $\mathcal{V}^{p}_{X}$ and $\mathcal{V}^{p}_{\tilde{X}}$ are, in general, different Banach spaces; if $X=\tilde{X}$, we write $\|\cdot\,;\cdot\|_{X,p,[s,t]}$.
@@ 2.2. Existence and uniqueness
To prove the existence of a unique solution to the rough functional differential equation (2.2), we postulate a quadratic growth and a Lipschitz type condition on the path-dependent coefficient $(F,F^{\prime})$, formulated on the associated path spaces. While a Lipschitz type condition is expected, the quadratic growth condition appears to be natural in the presented context of (second order) controlled paths, which corresponds to a Taylor expansion with quadratic remainder term.
@@
For any $X,\tilde{X}\in D^{p}([0,T];\mathbb{R}^{d})$ and for every $K>0$, there exist constants $C_{F}>0$, which depends on $p$ and the functional $F$, and $C_{F,K,X,\tilde{X}}>0$, which additionally depends on $K$, $X$, $\tilde{X}$ such that the maps
$(F,F^{\prime})\colon\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$
satisfy, for all $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}$, and every $0\leq s<t\leq T$,
@@
$\|F(Y)-F(\tilde{Y})\|_{p,[s,t]}\leq C_{F,K,X,\tilde{X}}(|Y_{s}-\tilde{Y}_{s}|+\|Y-\tilde{Y}\|_{p,[s,t]}),\quad\text{and}$
$\|F(Y),F^{\prime}(Y,Y^{\prime});F(\tilde{Y}),F^{\prime}(\tilde{Y},\tilde{Y}^{\prime})\|_{X,\tilde{X},p,[s,t]}$
$\quad\leq C_{F,K,X,\tilde{X}}(|Y_{s}-\tilde{Y}_{s}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[s,t]}+\|X-\tilde{X}\|_{p,[s,t]}),$
if $\|Y,Y^{\prime}\|_{X,p,[s,t]},\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p,[s,t]}\leq K$.
Remark 2.2.
When refering to a functional $(F,F^{\prime})$ which satisfies Assumption 2.1, we consider a functional which maps a controlled path with respect to some path $X\in D^{p}$ to a controlled path with respect to $X$, where $X$ is not being fixed a priori. That is, $(F(Y),F^{\prime}(Y,Y^{\prime}))\in\mathcal{V}^{p}_{X}$ for $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, for any $X\in D^{p}$. Particularly, Assumption 2.1 is to hold for any $X,\tilde{X}\in D^{p}$, which are not assumed fixed.
@@
Let $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ be a càdlàg $p$-rough path for $p\in(2,3)$, and $(y,y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be a given controlled path with respect to $X$. Suppose that the non-anticipative functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ satisfies Assumption 2.1. Then, there exists a unique solution to the rough functional differential equation (2.2), i.e. there exists a unique controlled path $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, with $Y^{\prime}=y^{\prime}+F(Y)$, such that
@@
$\mathcal{B}_{t}:=\bigg{\{}(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,t];\mathbb{R}^{k})\,:\,\begin{array}[]{l}(Y_{0},Y_{0}^{\prime})=(y_{0},y^{\prime}_{0}+F_{0}(y)),\\ \|(Y-y)^{\prime}\|_{p,[0,t]}\leq 4C_{F},\|R^{Y}-R^{y}\|_{\frac{p}{2},[0,t]}\leq 1\end{array}\bigg{\}}.$
@@
Contraction. Let $(Y,Y^{\prime}),(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{B}_{t}$ for some $t\in(0,t_{1}]$, that is, setting $K:=5(1+\|y,y^{\prime}\|_{X,p}+C_{F})$, it holds that $\|Y,Y^{\prime}\|_{X,p,[0,t]},\|\tilde{Y},\tilde{Y}^{\prime}\|_{p,[0,t]}\leq K$. We have that
$\|Z^{\prime}-\tilde{Z}^{\prime}\|_{p,[0,t]}=\|F(Y)-F(\tilde{Y})\|_{p,[0,t]}$
$\quad\leq C_{F,K,X,X}\|Y-\tilde{Y}\|_{p,[0,t]}$
$\quad\lesssim_{p}C_{F,K,X,X}(\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}\|X\|_{p,[0,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]}).$
@@
$\|R^{Z}-R^{\tilde{Z}}\|_{\frac{p}{2},[0,t]}=\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}F(\tilde{Y})\mathrm{d}\mathbf{X}}\|_{\frac{p}{2},[0,t]}$
$\quad\lesssim C_{F,K,X,X}\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,p,[0,t]}\|\mathbf{X}\|_{p,[0,t]},$
@@
$\|Z-\tilde{Z},Z^{\prime}-\tilde{Z}^{\prime}\|^{(\delta)}_{X,p,[0,t]}$
$\quad\lesssim C_{F,K,X,X}(\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}\|X\|_{p,[0,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]})$
$\qquad+\delta C_{F,K,X,X}(\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]})\|\mathbf{X}\|_{p,[0,t]}$
$\quad\lesssim C_{F,K,X,X}(1+\delta)\|\mathbf{X}\|_{p,[0,t]}\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}+C_{F,K,X,X}(1+\delta\|\mathbf{X}\|_{p,[0,t]})\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]}$
$\quad\lesssim C_{F,K,X,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t]}}{\delta}\bigg{)}\|Y-\tilde{Y},Y^{\prime}-\tilde{Y}^{\prime}\|^{(\delta)}_{X,p,[0,t]},$
@@
$C_{F,K,X,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t_{2}]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t_{2}]}}{\delta}\bigg{)}\leq 1,$
@@
We infer from Step 1 that there exists a constant $\gamma>0$, which depends on $p$, $\|y,y^{\prime}\|_{X,p}$, $C_{F}$, $C_{F,K,X,X}$ and $\|\mathbf{X}\|_{p}$, such that the local solution $(Y,Y^{\prime})$ established above exists on any interval $[s,t]$ such that $w(s,t)\leq\gamma$, given any initial condition $\xi\in\mathcal{V}^{p}_{X}$ with $|\xi^{\prime}_{s}|\leq\|y,y^{\prime}\|_{X,p}$.
@@
As stated in Step 2, we obtain the existence of a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\dots<\tau_{N}=T\}$ and define the solution on $(Y,Y^{\prime})$ on each of the half-intervals $[t_{i},t_{i+1})$, which then satisfies
@@
for all $i=0,\ldots,N-1$. Here, $N$ depends on $p$, $\|y,y^{\prime}\|_{X,p}$, $C_{F}$, $C_{F,K,X,X}$, $\|\mathbf{X}\|_{p}$ and is, for $p$ fixed, non-decreasing in the other variables. Observe that
@@
(2.5) $\|R^{Y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}\leq 1+\|R^{y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}+Q_{p}^{(R)}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
@@
(2.6) $\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1}]}\leq Q_{p}^{(Y^{\prime})}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]}).$

### v3

@@ 1. Introduction
where the driving signal $\mathbf{X}$ is a càdlàg $p$-rough path for $p\in(2,3)$, the initial condition $y$ is a given controlled path, and the coefficients $b,\sigma$ are non-anticipative functionals, mapping a controlled path to a controlled path. Assuming a quadratic growth and a Lipschitz-type condition on the path-dependent coefficients $b,\sigma$, which both are formulated on the space of controlled paths, we establish the existence of a unique solution to the RFDE (2.2). To that end, we rely on the theory of càdlàg rough paths, as introduced by Friz, Shekhar and Zhang [FS17, FZ18], as well as Banach’s fixed point theorem. Moreover, we show that the solution map, also known as Itô–Lyons map, mapping the input (initial condition, coefficients, driving signal) of an RFDE to its solution, is locally Lipschitz continuous with respect to suitable distances on the associated spaces of coefficients, controlled paths and rough paths. Let us remark that the continuity of the Itô–Lyons map is one of the most fundamental insights of rough path theory, with many applications to stochastic differential equations, cf. e.g. [FH20].
@@
In the existing literature, there are several different approaches to deal with rough functional differential equations driven by continuous rough paths. Since the theory of (continuous) rough paths works nicely for infinite dimensional Banach spaces, RDEs with path-dependent coefficients can be treated as Banach space-valued RDEs, see e.g. [Bai14], which requires the coefficients to be Fréchet differentiable and, thus, excludes some interesting examples. Existence, uniqueness and stability results are established by Neuenkirch, Nourdin and Tindel [NNT08] for RDEs with constant delay. The existence of a solution is proven by Ananova [Ana23] for RDEs with path-dependent coefficients, which are assumed to be Dupire differentiable [Dup19], and by Aida [Aid24] for RDEs with coefficients containing path-dependent bounded variation terms. The latter two approaches rely on Schauder’s fixed point theorem. Another exemplary class of RFDEs are reflected rough differential equations, see e.g. [Aid15, DGHT19], which, in general, do not possess a unique solution, see [Gas21].
@@
Most applications of rough path theory to stochastic differential equations (SDEs) crucially rely on the construction of suitable (random) rough paths. To apply the developed theory on RFDEs to Itô SDEs with constant delay, we show that a càdlàg martingale together with its delayed version can be lifted to a random rough path in the spirit of stochastic Itô integration. The key challenge to obtain the “delayed” rough path is that a martingale together with its delayed version is, in general, not a martingale itself, thus preventing the direct use of stochastic Itô integration. For related constructions of random rough paths above fractional Brownian motions we refer to [NNT08, TT12, BMCR14, CE20]. Consequently, one can apply the continuity of the Itô–Lyons map to derive pathwise stability results for stochastic differential equations with constant delay, which plays an important role in many applications, see e.g. [BDNHP18]. In particular, the map $y\mapsto Y$, mapping the initial condition $y$ to the associated solution $Y$ of an SDE with constant delay, is continuous on the space of controlled paths. This resolves an old observation, pointed out by Mohammed [Moh86], about the non-continuity of the flow of stochastic differential equations with delay, for which the initial condition is in fact an initial path.
@@
Organization of the paper: In Section 2 we provide existence, uniqueness and continuity results for rough functional differential equations. In Section 3 we prove that various classes of rough differential equations are covered by the presented results on rough functional differential equations. In Section 4 we establish the existence of the Itô rough path lift of delayed martingales and discuss applications to stochastic differential equations with delay. Appendix A contains some auxiliary estimates for rough integrals.
Acknowledgments: A. P. Kwossek and D. J. Prömel gratefully acknowledge financial support by the Baden-Württemberg Stiftung. A. P. Kwossek was affiliated with the University of Mannheim for the majority of this project’s duration.
@@ 2.1. Essentials on rough path theory
For two vector spaces, the space of linear maps from $E_{1}\to E_{2}$ is denoted by $\mathcal{L}(E_{1};E_{2})$; and we write $C^{l}_{b}=C^{l}_{b}(\mathbb{R}^{m};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ for the space of $l$-times differentiable (in the Fréchet sense) functions $f\colon\mathbb{R}^{m}\to\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})$ such that $f$ and all its derivatives up to order $l$ are continuous and bounded. We equip this space with the norm
@@
$X_{s,t}:=X_{t}-X_{s}\quad\text{and}\quad X_{t-}:=\lim_{u\uparrow t}X_{u},\qquad\text{for}\quad(s,t)\in\Delta_{T}.$
@@
$\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p}:=\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[0,T]}\qquad\text{with}\qquad\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[s,t]}:=\|X-\widetilde{X}\|_{p,[s,t]}+\|\mathbb{X}-\widetilde{\mathbb{X}}\|_{\frac{p}{2},[s,t]},$
@@
$\|Y\|_{p}\leq C_{p}(\|Y^{\prime}\|_{\infty}\|X\|_{p}+\|R^{Y}\|_{\frac{p}{2}})\qquad\text{and}\qquad\|Y^{\prime}\|_{\infty}\leq|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p},$
where $\|Y^{\prime}\|_{\infty}:=\sup_{t\in[0,T]}|Y^{\prime}_{t}|$ denotes the supremum seminorm of the path $Y^{\prime}$.
Given $X,\widetilde{X}\in D^{p}$, we further introduce the standard “distance”
$\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p}:=\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p,[0,T]}$
@@
$\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p,[s,t]}:=|Y^{\prime}_{s}-\widetilde{Y}^{\prime}_{s}|+\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[s,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[s,t]}$
for $(s,t)\in\Delta_{T}$, whenever $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X},(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}$. Note that $\mathcal{V}^{p}_{X}$ and $\mathcal{V}^{p}_{\widetilde{X}}$ are, in general, different Banach spaces; if $X=\widetilde{X}$, we write $\|\cdot\,;\cdot\|_{X,p,[s,t]}$.
@@ 2.2. Existence and uniqueness
To prove the existence of a unique solution to the rough functional differential equation (2.2), we postulate a quadratic growth and a Lipschitz-type condition on the path-dependent coefficient $(F,F^{\prime})$, formulated on the associated path spaces. While a Lipschitz-type condition is expected, the quadratic growth condition appears to be natural in the presented context of (second order) controlled paths, which corresponds to a Taylor expansion with quadratic remainder term.
@@
Let $X\in D^{p}([0,T];\mathbb{R}^{d})$ be given. For every $K>0$, there exist constants $C_{F}>0$, which depends on $p$ and the functional $F$, and $C_{F,K,X}>0$, which additionally depends on $K$ and $X$, such that the map
$(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$
satisfy, for all $(Y,Y^{\prime}),(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}$, and every $0\leq s<t\leq T$,
@@
$\|F(Y)-F(\widetilde{Y})\|_{p,[s,t]}\leq C_{F,K,X}(|Y_{s}-\widetilde{Y}_{s}|+\|Y-\widetilde{Y}\|_{p,[s,t]}),\quad\text{and}$
$\|F(Y),F^{\prime}(Y,Y^{\prime});F(\widetilde{Y}),F^{\prime}(\widetilde{Y},\widetilde{Y}^{\prime})\|_{X,p,[s,t]}$
$\quad\leq C_{F,K,X}(|Y_{s}-\widetilde{Y}_{s}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p,[s,t]}),$
if $\|Y,Y^{\prime}\|_{X,p,[s,t]},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p,[s,t]}\leq K$.
@@
Let $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ be a càdlàg $p$-rough path for $p\in(2,3)$, and $(y,y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be a given controlled path with respect to $X$. Suppose that the non-anticipative functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ satisfies Assumption 2.1 given $X$. Then, there exists a unique solution to the rough functional differential equation (2.2), i.e. there exists a unique controlled path $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, with $Y^{\prime}=y^{\prime}+F(Y)$, such that
@@
$\mathcal{B}_{t}:=\bigg{\{}(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,t];\mathbb{R}^{k})\,:\,\begin{array}[]{l}(Y_{0},Y_{0}^{\prime})=(y_{0},y^{\prime}_{0}+F_{0}(y)),\\ \|(Y-y)^{\prime}\|_{p,[0,t]}\leq 4C_{F},\|R^{Y}-R^{y}\|_{\frac{p}{2},[0,t]}\leq 1\end{array}\bigg{\}},$
which is a complete set as a closed subset of $\mathcal{V}^{p}_{X}([0,t];\mathbb{R}^{k})$, cf. [FZ18, Section 3.2].
@@
Contraction. Let $(Y,Y^{\prime}),(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{B}_{t}$ for some $t\in(0,t_{1}]$, that is, setting $K:=5(1+\|y,y^{\prime}\|_{X,p}+C_{F})$, it holds that $\|Y,Y^{\prime}\|_{X,p,[0,t]},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p,[0,t]}\leq K$. We have that
$\|Z^{\prime}-\widetilde{Z}^{\prime}\|_{p,[0,t]}=\|F(Y)-F(\widetilde{Y})\|_{p,[0,t]}$
$\quad\leq C_{F,K,X}\|Y-\widetilde{Y}\|_{p,[0,t]}$
$\quad\lesssim_{p}C_{F,K,X}(\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}\|X\|_{p,[0,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]}).$
@@
$\|R^{Z}-R^{\widetilde{Z}}\|_{\frac{p}{2},[0,t]}=\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}F(\widetilde{Y})\mathrm{d}\mathbf{X}}\|_{\frac{p}{2},[0,t]}$
$\quad\lesssim C_{F,K,X}\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p,[0,t]}\|\mathbf{X}\|_{p,[0,t]},$
@@
$\|Z-\widetilde{Z},Z^{\prime}-\widetilde{Z}^{\prime}\|^{(\delta)}_{X,p,[0,t]}$
$\quad\lesssim C_{F,K,X}(\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}\|X\|_{p,[0,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]})$
$\qquad+\delta C_{F,K,X}(\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]})\|\mathbf{X}\|_{p,[0,t]}$
$\quad\lesssim C_{F,K,X}(1+\delta)\|\mathbf{X}\|_{p,[0,t]}\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}+C_{F,K,X}(1+\delta\|\mathbf{X}\|_{p,[0,t]})\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]}$
$\quad\lesssim C_{F,K,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t]}}{\delta}\bigg{)}\|Y-\widetilde{Y},Y^{\prime}-\widetilde{Y}^{\prime}\|^{(\delta)}_{X,p,[0,t]},$
@@
$C_{F,K,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t_{2}]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t_{2}]}}{\delta}\bigg{)}\leq 1,$
@@
We infer from Step 1 that there exists a constant $\gamma>0$, which depends on $p$, $\|y,y^{\prime}\|_{X,p}$, $C_{F}$, $C_{F,K,X}$ and $\|\mathbf{X}\|_{p}$, such that the local solution $(Y,Y^{\prime})$ established above exists on any interval $[s,t]$ such that $w(s,t)\leq\gamma$, given any initial condition $\xi\in\mathcal{V}^{p}_{X}$ with $|\xi^{\prime}_{s}|\leq\|y,y^{\prime}\|_{X,p}$.
@@
Analogously to Step 2, we can obtain a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\dots<\tau_{N}=T\}$ and, recalling the definition of $\mathcal{B}_{t}$, define the solution $(Y,Y^{\prime})$ on each of the half-intervals $[t_{i},t_{i+1})$ such that
@@
for all $i=0,\ldots,N-1$. Here, $N$ depends on $p$, $\|y,y^{\prime}\|_{X,p}$, $C_{F}$, $C_{F,K,X}$, $\|\mathbf{X}\|_{p}$ and is, for $p$ fixed, non-decreasing in the other variables. Observe that
@@
$\|R^{Y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}\leq 1+\|R^{y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}+Q_{p}^{(R)}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
@@
$\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1}]}\leq Q_{p}^{(Y^{\prime})}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]}).$